bruhhurb:

Identify the variation as direct, inverse, joint or combined. xy = 24

tiaunabuchan:

what do u think

jhonyy9:

so for example if \[\frac{ x }{ y } = 24 \] is inverse variation how you think that will be x * y = 24 ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XxAngelGirlxX xy = 24 is an inverse variation. (Please close when done) \(\color{#0cbb34}{\text{End of Quote}}\) no x/y = 24 is an inverse so x*y = 24 not can be too inverse